Session: Feb 1997 (session 1)
Frequency: 6 cm
Obs time (UT): EU 08:00(21/02)-19:00(21/02)
EU Mk3 Tapes/St.: 3
To give feedback on this experiment, click here
Westerbork (Tony Foley at Fri Feb 21 09:45:35 EST 1997) Minor failures
Lost first scan (0800-0803) as telescopes were
not on source. Experiment ended early at Westerbork
as the sources were outside our hour angle range
Medicina (D.Dallacasa at Fri Feb 21 14:48:49 EST 1997) Success
Source sets 45 min before end of the schedule
Onsala (Antonis Polatidis at Sun Feb 23 09:00:54 EST 1997) Minor failures
Due to high winds we had to stow the antenna at UT 1105.
Effelsberg (dgraham at Mon Feb 24 03:28:04 EST 1997) Success
OK. This was not supposed to be a double speed experiment.
Jodrell (Alastair Gunn at Wed Mar 5 08:36:37 EST 1997) Minor failures
No major problems. Missed scan before 16:41 due to azimuth
un-wrapping. Tapes EVN0086,CA90 IRAM0084,078F and JODV0062,C8B9
sent to Socorro via Fedex on 24/02/97.
Cambridge (Alastair Gunn at Wed Mar 5 08:44:00 EST 1997) Minor failures
No major problems. Formatter clock discovered to be behind GPS
by 2 seconds and fixed for 09:40 scan. Problem was not present
one hour before this. Tapes EVN00118,C6FC MPIR1184,B747 and
EVN00035,5ACD sent to Socorro via Fedex on 24/02/97.
Calibration (Fredrik T Rantakyro at Thu May 8 11:34:49 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the
Feb97 directory with the name gp15a.antab. This file contains the
preliminary calibration results and may be updated when I receive
updated information from the telescopes and the results from the
datareduction of the calibration experiments.